OpenStudy (julietstar1):

How do you factor 2y^3 + 8y^2 + 8y?

OpenStudy (anonymous):

well you can factor the whole polynomial by 2y 2y(y^2+4y+4) Then factor the remaining polynomial (2y)(y+2)(y+2)

OpenStudy (psymon):

Find the largest number that can evenly be divided out of all the terms. Then find the highst power of y that can be divided out of all terms. Divide those values out and rewrite it like a distributive property. For example: 3x^3 + 6x^2 + 9x = 3x(x^2 + 2x + 3) I literally divided all terms by the highest common factor, and then pulled that out in front and rewrote it. Then you factor what is remaining afterwards. But it was done for you, so yeah.

OpenStudy (anonymous):

Since they all have a "y" and a "2", factor those out first to give you ==> 2y ( y^2 + 4y + 4) We can then factor the polynomial in parenthesis y^2 + 4y + 4 = (y+2)(y+2) = (y+2)^2 plug that into the original equation to give you ==> \[2y (y+2)^{2}\]